# Learning User Embeddings from Human Gaze for Personalised Saliency Prediction

### v1

@@ 3.2. Personalised Saliency Prediction
[…]
@@ 4.1. Implementation Details
Model
CC
SIM
AUC
NSS
KLD
DeepGazeIIE (DG)
0.622
0.556
0.904
2.121
1.123
Fine-tuned DG
0.715
0.619
0.905
2.140
0.526
MultiCNN w/ DG 0.735 0.643
0.897
2.142
0.708
Ours w/ DG 0.736 $*{}^{*}$ 0.907 $*{}^{*}$ $*{}^{*}$
Ground Truth USM
0.801
0.685 0.921
2.373
0.372
MultiCNN w/ GT 0.804
0.683
0.919
2.378
0.511
Ours w/ GT $*{}^{*}$ $*{}^{*}$ 0.922 $*{}^{*}$ $*{}^{*}$
@@
Model
CC
SIM
AUC
NSS
KLD
DeepGazeIIE (DG)
0.584
0.530
0.881
2.288
1.294
Fine-tuned DG
0.734
0.622
0.887
2.293
0.562
MultiCNN w/ DG 0.746 0.637 0.892 2.299
0.604
Ours w/ DG $*{}^{*}$ $*{}^{*}$ 0.896 2.308 $*{}^{*}$
Ground Truth USM
0.821
0.698
0.912
2.500
0.366
MultiCNN w/ GT 0.845 0.711 0.914 2.609
0.415
Ours w/ GT 0.846 $*{}^{*}$ 0.915 2.595 $*{}^{*}$
@@ 4.3. Baselines
Model
CC
SIM
AUC
NSS
KLD
DeepGazeIIE (DG)
0.615
0.553
0.916
2.239
1.083
Fine-tuned DG
0.721
0.624
0.914
2.244
0.542
MultiCNN w/ DG 0.723 0.627
0.910
2.244
0.620
Ours w/ DG 0.726 $*{}^{*}$ 0.916 2.253 $*{}^{*}$
Ground Truth USM
0.804
0.682
0.922
2.506
0.387
MultiCNN w/ GT 0.804 0.690 0.929 2.550
0.440
Ours w/ GT $*{}^{*}$ $*{}^{*}$ 0.933 2.552 $*{}^{*}$
@@
Model
CC
SIM
AUC
NSS
KLD
DeepGazeIIE (DG)
0.564
0.522
0.860
1.944
1.431
Fine-tuned DG
0.692
0.618
0.862
2.032
0.685
MultiCNN w/ DG 0.693
0.615
0.862
2.006
0.682
Ours w/ DG
0.712
0.634
0.870
2.033
0.539
Ours CD w/ DG $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ 2.033 $*{}^{*}$
Ground Truth USM
0.788
0.691
0.892
2.218
0.399
MultiCNN w/ GT 0.800 0.694 0.892 2.290
0.516
Ours w/ GT
0.804
0.701
0.893
2.273
0.371
Ours CD w/ GT 0.807 $*{}^{*}$ 0.894 2.288 $*{}^{*}$

### v2

@@ 3.2. Personalised Saliency Prediction
where $N(...)_{k}$ is one of the three predicted discrepancy maps of the personalised saliency network. To calculate the ground truth discrepancy maps $\Delta(I,U)$ we use Equation 2 and subtract the USMI from the PSMI,U.
@@ 4.1. Implementation Details
Model CC SIM AUC NSS KLD
DeepGazeIIE (DG) 0.622 0.556 0.904 2.121 1.123
Fine-tuned DG 0.715 0.619 0.905 2.140 0.526
MultiCNN w/ DG 0.735 0.643 0.897 2.142 0.708
Ours w/ DG 0.736 0.651 0.907 2.170 0.509
Ground Truth USM 0.801 0.685 0.921 2.373 0.372
MultiCNN w/ GT 0.804 0.683 0.919 2.378 0.511
Ours w/ GT 0.813 0.706 0.922 2.405 0.357
@@
Model CC SIM AUC NSS KLD
DeepGazeIIE (DG) 0.584 0.530 0.881 2.288 1.294
Fine-tuned DG 0.734 0.622 0.887 2.293 0.562
MultiCNN w/ DG 0.746 0.637 0.892 2.299 0.604
Ours w/ DG 0.760 0.649 0.896 2.308 0.481
Ground Truth USM 0.821 0.698 0.912 2.500 0.366
MultiCNN w/ GT 0.845 0.711 0.914 2.609 0.415
Ours w/ GT 0.846 0.725 0.915 2.595 0.336
@@ 4.3. Baselines
Model CC SIM AUC NSS KLD
DeepGazeIIE (DG) 0.615 0.553 0.916 2.239 1.083
Fine-tuned DG 0.721 0.624 0.914 2.244 0.542
MultiCNN w/ DG 0.723 0.627 0.910 2.244 0.620
Ours w/ DG 0.726 0.638 0.916 2.253 0.527
Ground Truth USM 0.804 0.682 0.922 2.506 0.387
MultiCNN w/ GT 0.804 0.690 0.929 2.550 0.440
Ours w/ GT 0.813 0.702 0.933 2.552 0.366
@@
Model CC SIM AUC NSS KLD
DeepGazeIIE (DG) 0.564 0.522 0.860 1.944 1.431
Fine-tuned DG 0.692 0.618 0.862 2.032 0.685
MultiCNN w/ DG 0.693 0.615 0.862 2.006 0.682
Ours w/ DG 0.712 0.634 0.870 2.033 0.539
Ours CD w/ DG 0.713 0.634 0.871 2.033 0.533
Ground Truth USM 0.788 0.691 0.892 2.218 0.399
MultiCNN w/ GT 0.800 0.694 0.892 2.290 0.516
Ours w/ GT 0.804 0.701 0.893 2.273 0.371
Ours CD w/ GT 0.807 0.704 0.894 2.288 0.361
@@ 6. Summary
Acknowledgements.
Florian Strohm and Andreas Bulling were funded by the European Research Council (ERC) under the grant agreement 801708. Mihai Bâce was funded by a Swiss National Science Foundation (SNSF) Postdoc.Mobility Fellowship (grant number 214434).